Please help,thank you.
The length of a rectangle is 1 cm longer than it's width. If the diagonal of the rectangle is 4cm,what are the dimensions(the length and width)of the rectangle?

Answer by checkley75(3666)

LENGTH=W+1 & WIDTH=W
THUS THE DIAGONAL OR HYPOTENUSE (SQUARED 4^2)=L^2+W^2
L^2+W^2=4^2
(W+1)^2+W^2=16
W^2+2W+1+W^2=16
2W^2+2W+1-16=0
2W^2+2W-15=0
USING THE QUADRATIC EQUATION WE GET
X=(-2+-SQRT[2^2-4*2*-15])/2*2
X=(-2+-SQRT[4+120])/4
X=(-2+-SQRT[124])/4
X=(-2+-11.13)/4
X=(-2+11.13)/4
X=9.13/4
X=2.28 ANSWER
X=(-2-11.13)/4
X=-13.13/4
X=-3.28 ANSWER